If demand P=100-Q, TC=800+20Q+Q2 and MC=20+2Q what is the profit maximizing price and the rate of output


1
Expert's answer
2021-08-29T16:53:21-0400

profit=total revenue(TR)-total cost(TC)

Given,

"TC = 800+20Q+Q^2\\

P = 100 - Q\\

TR = Price (P) \times Quantity (Q)\\

TR = (100 - Q)\times Q\\

TR = 100Q - Q^2\\

So\space here,\\ 

Profit = 100Q - Q^2 - (800+20Q+Q^2)\\

 = 100Q - Q^2 - 800 - 20Q - Q^2\\

 = 80Q -2Q^2 - 800"

Now, for maximizing profit we will derivate profit with respect to Q and will put it equal to zero

"\frac{∂\space Profit }{ ∂Q} = 80 - 4Q\\

Putting\\

\frac{∂Profit }{ ∂Q} = 0\\ 

80 - 4Q = 0\\

80 = 4Q\\

Q = 20"


So, rate of output is 20

Putting value of Q in P


"P = 100 - 20\\

P = 80"


So, profit maximizing price is 80 and rate of output is 20.